. Paul has $160 worth of $20 bills. He has half as many 10s as he does 20s. He has half as many 5s as he does 10s. And he has half as many 1s as he does 5s. How much money does Paul have?
Answer by macston(5194)

Paul has $160 worth of $20 bills. He has half as many 10s as he does 20s. He has half as many 5s as he does 10s. And he has half as many 1s as he does 5s. How much money does Paul have?
Tw=number of twenties; T=number of tens; F=number of fives; N=number of ones
$160/$20=8 Paul has 8 $20 bills.
T=(1/2)Tw=4 Paul has 4 $10 bills.
F=(1/2)T=2 Paul has 2 $5 bills.
N=(1/2)F=1 Paul has 1 $1 bill.
Total money=$20Tw+$10T+$5F+$1N
Total money=$20(8)+$10(4)+$5(2)+$1(1)
Total money=$160+$40+$10+$1=$211
ANSWER: Paul has $211.